OpenStudy (anonymous):

Convert the following to rectangular form: 2(cos(135)+isin(135)) 3(cos(120)+isin(120)) 5(cos(5π/4)+isin(5π/4)) 4(cos(5π/3)+isin(5π/3))

OpenStudy (unklerhaukus):

\[re^{i\theta} = r(\cos\theta+i\sin\theta)=r(\cos\theta,\, \sin\theta)=(r\cos\theta,\, r\sin\theta)\]

OpenStudy (cwrw238):

are you clear on that 11SAAT?